# Hydrodynamic memory and Quincke rotation

### v1

@@ II.1 Lorenz equations with simplified memory (sMLE)
In above, we have assumed that $X(0)=0$  [^1] Even if $X(0)\neq 0$, the additional term $X(0)M(s)$ goes to zero exponential as $s\xrightarrow{}\infty$.. Now if we denote the history integral term by $H(s)=\int_{-\infty}^{s}X(s^{\prime})M(s-s^{\prime})\,\mathrm{d}s^{\prime}$ then we get the system of equations
@@
III Linear Stability Analysis of a stationary particle
III.1 MLE model
The MLE has a trivial solution of $(X,Y,Z)=(0,0,0)$, corresponding to a stationary particle whose dipole is anti-parallel to the applied field. We examine the stability of this base state by introducing a small perturbation $(X,Y,Z)=(0,0,0)+\delta(X_{1},Y_{1},Z_{1})$, where $\delta\ll 1$, into the MLE and neglecting any terms quadratic or higher powers in $\delta$. This yields the following initial value problem for $(X_{1},Y_{1},Z_{1})$,
@@ III.1 MLE model
where an overhead ()similar-to($\sim$)( ∼ ) denotes the Laplace transform, $\hat{s}$ is the Laplace frequency, and $\tilde{M}(\hat{s})=\gamma/(3[\sqrt{\gamma\hat{s}}+1])$. While we are unable to invert (19) in closed form, progress is made at long times by taking the limit as $\hat{s}\rightarrow 0$ of $\hat{s}\tilde{X}_{1}(\hat{s})$, thereby yielding
$\lim_{\hat{s}\to 0}\hat{s}\tilde{X}_{1}(\hat{s})=-\frac{Pr\left(\frac{1}{3}\gamma X_{1}(0)-Y_{1}(0)\right)+X_{1}(0)}{Pr\left(r-1\right)}.$ (20)
Hence, for small $\hat{s}$, corresponding to long times, inversion of (20) yields exponential solutions $X_{1}(t)\sim\exp[3Pr(r-1)s]$. The initial perturbation grows (decays) if the argument of the exponential is positive (negative); the criterion for marginal stability is thus $r=1$, which in dimensional terms is $E_{a}=E_{c}$. A linear stability analysis of the standard Lorenz equations (LE) and the sMLE model (as shown below) also predicts the same criteria of instability of the stationary state. Therefore, hydrodynamic memory does not alter the threshold field strength for linear instability. Intuitively, this is expected since the base state is stationary.
III.2 sMLE model
@@
IV Linear Stability Analysis of steady rotation
IV.1 MLE model
Once the stationary state becomes unstable for $r>1$, the system transitions to another steady state $(X,Y,Z)=(\pm\sqrt{r-1},\pm\sqrt{r-1},r-1)$ corresponding to a steady angular velocity with clockwise or counterclockwise rotation. This steady state is also the same for LE and sMLE (as shown below). Thus, memory effects do not have an influence in setting the threshold for the primary pitchfork bifurcation as well as the corresponding steady state the system settles into after the bifurcation. Again, this is intuitive since at long times the particle is rotating steadily and the memory of its approach to that steady state fades as the memory kernel decays at large times. We examine the stability of the counterclockwise rotating steady state (without loss of generality) by introducing a small perturbation $(X,Y,Z)=(\sqrt{r-1},\sqrt{r-1},r-1)+\delta(X_{1},Y_{1},Z_{1})$, where $\delta\ll 1$, into the MLE and neglecting any terms quadratic or higher in $\delta$. This yields the following initial value problem for $(X_{1},Y_{1},Z_{1})$,
@@ IV.1 MLE model
The growth rates of this linear stability problem corresponds to the poles of $\mathbf{X}(\hat{s})$. Hence finding the growth rates reduces to determining the roots of $\det(\mathsf{A}(\hat{s}))=0$. We solve this numerically by finding solutions in the two dimensional space formed by real and imaginary part of $\hat{s}$ that satisfy both real and imaginary parts of $\det(\mathsf{A}(\hat{s}))=0$. The instability of steady rotation state takes place when the real part of $\hat{s}$ (that solves $\det(\mathsf{A}(\hat{s}))=0$) changes sign from negative to positive.
In figure 2(a) the marginal stability curves for the steady rotation branch in $r-Pr$ space are plotted for the MLE. It can be seen that an increase in memory (i.e. increasing $\gamma$) leads to a stabilization of the steady rotation state in the parameter space compared to the standard Lorenz equations (LE) shown as a gray curve. That is the critical value of $r$ for instability of the steady rotation increases with $\gamma$ for a fixed $Pr$.
IV.2 sMLE model
@@
In figure 2(b) the marginal stability curves for the steady rotation branch in $r-Pr$ space are plotted for sMLE. It can be seen that here too, an increase in memory (i.e. increasing $\alpha$) leads to a stabilization of the steady rotation state in the parameter space compared to LE.
Figure 2: Linear stability curves for steady rotation. Curves showing the linear stability boundary of the steady rotation branch in the parameter space of $Pr$ and $r$ for (a) MLE and (b) sMLE. The limit $\gamma\xrightarrow{}0$ and $\alpha\xrightarrow{}0$ corresponds to the Lorenz equations (LE), shown as a gray curve. This gray curve corresponding to instability of the steady state in LE takes the analytical form $r=Pr(Pr+4)/(Pr-2)$ [20].
IV.3 Beyond steady rotation
For larger $r$ values beyond steady rotation, a complex set of bifurcations take place in LE [29, 30]. Furthermore, global bifurcations in LE take place even when the steady rotation state is stable resulting in coexistence of steady states and chaotic dynamics [29, 30]. Similar level of complexity would also be expected for sMLE and MLE. Nonlinear dynamics arising from global bifurcations cannot be captured with a linear analysis near fixed points, and one would need to resort to numerical simulations; we explore these in section VI. To investigate into these nonlinear behaviors, we would need to numerically solve the three models. It is straightforward to solve the ODEs corresponding to LE and sMLE and we do this in MATLAB using the inbuilt solver ode45. However numerically solving the integro-differential equations for MLE is not trivial. We describe this numerical procedure in the next section.
V Numerical Integration of MLE
@@
VI Nonlinear Quincke Rotor Dynamics
@@
Figure 3: (Left) Time series of the dimensionless angular velocity $X$ of the Quincke rotor for (blue) standard Lorenz system, (red) sMLE with $\alpha=0.7$, and (black) MLE with $\gamma=1$ at the dimensionless applied electric field strength (a) $r=3$, (b) $r=52$ and (c) $r=105$. (Right) Corresponding basin of attraction plots in the initial condition space formed by $(Y(0),Z(0))$ for sMLE model showing the final value of $X$ at the end of the simulation (with simulation time of $s=500$) for fixed $X(0)=0$. We see a smoothly separated basin for $r=3$, fractal basins due to transient chaos for $r=52$, and noise in the plot for $r=105$ due to permanent chaos. The parameter $Pr$ was fixed to $2.5$. Supplemental Videos 1-3 show the dynamics of MLE at $r=3,52$ and $105$, respectively.
@@
We first calculate the time-series dynamics of a spherical particle undergoing Quincke rotation with $\gamma=1$ in MLE for three values of $r$, the dimensionless applied electric field strength. We compare these results at the same $r$ value for sMLE (with $\alpha=0.7$) and LE models. In figure 3, the time series for all three models is shown for $r=3$, $52$, and $105$, demonstrating steady rotation ($r=3$); transient chaos (or ‘preturbulence’) before steady rotation ($r=52$); and permanent chaotic dynamics ($r=105)$. For $r=3$, we find that all three models, LE (blue), sMLE (red) and MLE (black), predict a steady rotation (see figure 3(a)). This steady state rotation value is given by $X=\pm\sqrt{r-1}$ from the steady equilibrium states of the three models. At this $r$ value, a basin of attraction plotting the final value of $X$ in the initial condition space formed by $Y(0)$ and $Z(0)$ using sMLE equations is shown in the right panel of figure 3(a). We see that the basin of clockwise and counterclockwise steady rotating states are smoothly partitioned.
@@
For a larger value of $r=52$, as shown in figure 3(b), we find that sMLE and MLE settles eventually on the steady value of $X$ after a chaotic transient, whereas LE, which does not include hydrodynamic memory, already exhibits permanent chaotic motion. Thus, we already see a signature of delay in the onset of chaotic motion with the inclusion of a memory term in the Lorenz equations. For LE at $Pr=2.5$, the steady rotation branch becomes unstable at $r=32.5$; hence, at $r=52$ the only attracting set in phase space is a strange attractor resulting in chaotic dynamics. We note that the presence of transient chaos for sMLE and MLE indicates a final state sensitivity based on the initial conditions. Unlike $r=3$, where a $Y(0)>0$ would always result in counterclockwise steady rotation and $Y(0)<0$ in clockwise steady rotation, now at $r=52$, whether a clockwise or a counterclockwise steady rotation will be achieved in sMLE and MLE depends sensitively on initial conditions due to transient chaos. This can be seen in the basin of attraction plot of sMLE in the right panel of figure 3(b), where the basins of the two steady states are intricately mixed with fractal boundaries; a characteristic feature of transient chaos [32].
@@
For an even larger value of $r=105$, as shown in figure 3(c), the sMLE and MLE systems have now transitioned to permanent chaos, whereas the LE shows a periodic trajectory. The chaotic nature of the trajectory results in a noisy basin plot for sMLE as shown in the right panel of figure 3(c), where the final velocity at the end of the simulation can now take any value and is not correlated with the initial conditions.
@@
Figure 4: Bifurcation diagram showing the peaks of $|X|$ i.e. $|X|_{n}$ as a function of $r$ and fixed $Pr=2.5$ for (a) standard LE, (b) sMLE with $M(s)=\alpha\,\text{e}^{-s/\alpha}$ and $\alpha=0.7$, and (c) MLE with $\gamma=1$. For each plot, the lower branch is the steady rotation state with the solid curve indicating where it is stable and the dashed curve indicating its unstable. The steady branch becomes unstable at $r=32.5$, $r\approx 92$ and $r\approx 176$ for LE, sMLE and MLE, respectively. Note the multistability regions where both chaotic motion and steady rotation coexist. The onset of multistability occurs at $r\approx 25$, $r\approx 59$ and $r\approx 55$ for LE, sMLE and MLE, respectively.
@@
In figure 4 we present a bifurcation diagram showing the peaks of $|X|$ as a function of $r$ and fixed $Pr=2.5$. Here the numerical integration was conducted until a time $s=500$ for the LE and sMLE, and $s=400$ for the MLE (due to the lengthier computations associated with numerical integration of the MLE). For the LE the steady rotation branch becomes unstable at $r=32.5$, but a multistability region where both chaotic motion and steady rotation coexist occurs below this value of $r$. Indeed, chaotic dynamics are first observed at $r\approx 25$. The onset of chaotic dynamics (coexisting with steady rotation) is seen at $r\approx 55$ for the MLE (with $\gamma=1)$, which is a significantly larger $r$ value than for the LE. This larger $r$ value for the second bifurcation in the MLE is also seen in experiments by Peters [20]; that is, the onset of chaos in their experiments occurs at a larger $r$ value than that predicted by the LE, and the present work suggests that hydrodynamic memory is the cause. The sMLE also have an increased $r$ value for the onset of chaotic dynamics (coexisiting with steady rotation) of $r\approx 59$,T which suggests this to be a generic consequence of hydrodynamic memory, irrespective of the precise form of the memory kernel.
@@
In figure 5 the multistability of the sMLE (with $\alpha=0.7$, figure 5(a)) and MLE (with $\gamma=1$, figure 5(b)) are examined by plotting the time series of $X$ for two different initial conditions, where one asymptotes towards a chaotic attractor (red) and the other towards a steady state (blue). Figure 5(c) displays basin of attraction plot in the initial condition space $(Y(0),Z(0))$ in the sMLE model for fixed $X(0)=0$ showing the final value of $X$ at the end of simulation time of $400$. We can see coexistence of steady states with smooth separation of their basins, which reside alongside chaotic solutions with noisy values in the basin. Thus, in the multistable region, we see that understanding the size and organization of basin of attraction of steady rotation and chaotic motion might assist with predicting which of the two final states is likely to be realized.
@@
Figure 5: Multistability at $Pr=2.5$ and $r=59$ in the (a) sMLE model with $\alpha=0.7$ and (b) MLE with $\gamma=1$. Time series of $X$ for two different initial conditions where one asymptotes towards a chaotic attractor (red) and the other towards a steady state (blue). (c) Corresponding basin of attraction plot in the initial condition space formed by $(Y(0),Z(0))$ for sMLE model showing the final value of $X$ for fixed $X(0)=0$. Supplemental Videos 4 and 5 show for MLE model the corresponding chaotic motion and steady dynamics, respectively.
@@
In figure 6(a) we present a bifurcation diagram of the MLE showing the peaks of $|X|$ as a function of $r$ at fixed $Pr=2.5$ and $\gamma=0.5$. Here, the onset of chaos is at $r\approx 42$ and there is again a region of multistability in which chaotic and steady motion coexist (with steady rotation stable up to $r\approx 90$). In figure 6(b) and figure 6(c) we plot the times series of $X$ just (b) before ($r=87$) and (c) after ($r=92$), respectively, the end of the stability of the steady rotation branch. At $r=87$ there is a slow converge of $X$ toward its steady value, and for $r=92$ there is a slow growth of $X$, which we envision would at some time point beyond the simulation become chaotic. The value of $\gamma=0.5$ was chosen as this is approximately the same value as in the experiments of [20]. They observed chaotic dynamics at a field strength of $6.5\mathrm{kV/cm}$, whereas the LE would predict chaos to onset at $5.5\mathrm{kV/cm}$, which corresponds to the ratio of $r$ values (experiment to LE) of around $1.4$. Our computations at $\gamma=0.5$ give a ratio of $r$ values (MLE to LE) of around $1.3$, which is in encouraging agreement with the experimental findings.
@@
In figure 7 the time series of $X(t)$ is plotted at a large value of $r=250$ for the LE, sMLE with $\alpha=0.7$, and MLE with $\gamma=1$. All three series converge onto a periodic orbit; however, the orbits are not in phase. It is well known for LE that periodic solutions exist at large $r$ and the system becomes integrable in this limit [30]. It appears that we get the same limiting behavior for sMLE and MLE for large $r$. It would be interesting to explore this regime experimentally and verify if periodic rotations are obtained for Quincke rotors under large magnitude of applied electric field strengths.
@@
Figure 6: (a) Bifurcation diagram showing the peaks of $|X|$ i.e. $|X|_{n}$ as a function of $r$ and fixed $Pr=2.5$ for full MLE with $\gamma=0.5$. The lower branch is the steady rotation state with the solid curve indicating where it is stable and the dashed curve indicating its unstable. (b) Time series of $X$ just (b) before ($r=87$) and (c) after ($r=92$) the end of the stability of the steady rotation branch.
@@
Figure 7: Time series of $X$ showing periodic oscillations for large $r=250$ and fixed $Pr=2.5$ for (a) LE, (b) sMLE with $\alpha=0.7$, and (c) full MLE with $\gamma=1$. Supplemental Video 6 shows the dynamics of MLE in this periodic state.
VII Conclusion

### v2

@@ III Lorenz equations with simplified memory (sMLE)
In above, we have assumed that $X(0)=0$  [^1] Even if $X(0)\neq 0$, the additional term $X(0)M(s)$ goes to zero exponential as $s\xrightarrow{}\infty$.. Now if we denote the history integral term by $H(s)=\int_{0}^{s}X(s^{\prime})M(s-s^{\prime})\,\mathrm{d}s^{\prime}$ then we get the system of equations
@@
IV Linear Stability Analysis of a stationary particle
IV.1 MLE model
The MLE has a trivial solution of $(X,Y,Z)=(0,0,0)$, corresponding to a stationary particle whose dipole is anti-parallel to the applied field. We examine the stability of this base state by introducing a small perturbation $(X,Y,Z)=(0,0,0)+\delta(X_{1},Y_{1},Z_{1})$, where $0<\delta\ll 1$, into the MLE and neglecting any terms quadratic or higher powers in $\delta$. This yields the following initial value problem for $(X_{1},Y_{1},Z_{1})$,
@@ IV.1 MLE model
[…]
$\lim_{\hat{s}\to 0}\tilde{X}_{1}(\hat{s})=\frac{Pr\left(\frac{1}{3}\gamma X_{1}(0)-Y_{1}(0)\right)+X_{1}(0)}{\hat{s}-Pr\left(r-1\right)}.$ (20)
Hence, for small $\hat{s}$, corresponding to long times, inversion of (20) yields exponential solutions $X_{1}(s)\sim\exp[3Pr(r-1)s]$. The initial perturbation grows (decays) if the argument of the exponential is positive (negative); the criterion for marginal stability is thus $r=1$, which in dimensional terms is $E_{a}=E_{c}$. A linear stability analysis of the standard Lorenz equations (LE) and the sMLE model (as shown below) also predicts the same criteria of instability of the stationary state. Therefore, hydrodynamic memory does not alter the threshold field strength for linear instability. Intuitively, this is expected since the base state is stationary.
IV.2 sMLE model
@@
V Linear Stability Analysis of steady rotation
V.1 MLE model
Once the stationary state becomes unstable for $r>1$, the system transitions to another steady state $(X,Y,Z)=(\pm\sqrt{r-1},\pm\sqrt{r-1},r-1)$ corresponding to a steady angular velocity with clockwise or counterclockwise rotation. This steady state is also the same for LE and sMLE (as shown below). Thus, memory effects do not have an influence in setting the threshold for the primary pitchfork bifurcation as well as the corresponding steady state the system settles into after the bifurcation. Again, this is intuitive since at long times the particle is rotating steadily and the memory of its approach to that steady state fades as the memory kernel decays at large times. We examine the stability of the counterclockwise rotating steady state (without loss of generality) by introducing a small perturbation $(X,Y,Z)=(\sqrt{r-1},\sqrt{r-1},r-1)+\delta(X_{1},Y_{1},Z_{1})$, where $0<\delta\ll 1$, into the MLE and neglecting any terms quadratic or higher in $\delta$. This yields the following initial value problem for $(X_{1},Y_{1},Z_{1})$,
@@ V.1 MLE model
[…]
In figure 2(a) the marginal stability curves for the steady rotation branch in rPrr-Pr space are plotted for the MLE. It can be seen that an increase in memory (i.e. increasing $\gamma$) leads to a stabilization of the steady rotation state in the parameter space compared to the standard Lorenz equations (LE) shown as a gray curve. That is the critical value of $r$ for instability of the steady rotation increases with $\gamma$ for a fixed $Pr$.
V.2 sMLE model
@@
In figure 2(b) the marginal stability curves for the steady rotation branch in rPrr-Pr space are plotted for sMLE. It can be seen that here too, an increase in memory (i.e. increasing $\alpha$) leads to a stabilization of the steady rotation state in the parameter space compared to LE.
Figure 2: Linear stability curves for steady rotation. Curves showing the linear stability boundary of the steady rotation branch in the parameter space of $Pr$ and $r$ for (a) MLE and (b) sMLE. The limit $\gamma\xrightarrow{}0$ and $\alpha\xrightarrow{}0$ corresponds to the Lorenz equations (LE), shown as a gray curve. This gray curve corresponding to instability of the steady state in LE takes the analytical form r=Pr(Pr+4)/(Pr2)r=Pr(Pr+4)/(Pr-2) [20].
V.3 Beyond steady rotation
For larger $r$ values beyond steady rotation, a complex set of bifurcations take place in LE [29, 30]. Furthermore, global bifurcations in LE take place even when the steady rotation state is stable resulting in coexistence of steady states and chaotic dynamics [29, 30]. Similar level of complexity would also be expected for sMLE and MLE. Nonlinear dynamics arising from global bifurcations cannot be captured with a linear analysis near fixed points, and one would need to resort to numerical simulations; we explore these in section VII. To investigate into these nonlinear behaviors, we would need to numerically solve the three models. It is straightforward to solve the ODEs corresponding to LE and sMLE and we do this in MATLAB using the inbuilt solver ode45. However numerically solving the integro-differential equations for MLE is not trivial. We describe this numerical procedure in the next section.
VI Numerical Integration of MLE
@@
VII Nonlinear Quincke Rotor Dynamics
@@
Figure 3: (Left) Time series of the dimensionless angular velocity $X$ of the Quincke rotor for (blue) standard Lorenz system, (red) sMLE with $\alpha=0.5$, and (black) MLE with $\gamma=1$ at the dimensionless applied electric field strength (a) $r=3$, (b) $r=47$ and (c) $r=110$. (Right) Corresponding basin of attraction plots in the initial condition space formed by $(Y(0),Z(0))$ for sMLE model showing the final value of $X$ at the end of the simulation (with simulation time of $s=500$) for fixed $X(0)=0$. We see a smoothly separated basin for $r=3$, fractal basins due to transient chaos for $r=47$, and noise in the plot for $r=110$ due to permanent chaos. The parameter PrPr was fixed to $2.5$. Supplemental Videos 1-3 show the dynamics of MLE at $r=3,47$ and $110$, respectively.
@@
We first calculate the time-series dynamics of a spherical particle undergoing Quincke rotation with $\gamma=1$ in MLE for three values of $r$, the dimensionless applied electric field strength. We compare these results at the same $r$ value for sMLE (with $\alpha=0.5$) and LE models. In figure 3, the time series for all three models is shown for $r=3$, $47$, and $110$, demonstrating steady rotation ($r=3$); transient chaos (or ‘preturbulence’) before steady rotation ($r=47$); and permanent chaotic dynamics ($r=110)$. For $r=3$, we find that all three models, LE (blue), sMLE (red) and MLE (black), predict a steady rotation (see figure 3(a)). This steady state rotation value is given by $X=\pm\sqrt{r-1}$ from the steady equilibrium states of the three models. At this $r$ value, a basin of attraction plotting the final value of $X$ in the initial condition space formed by $Y(0)$ and $Z(0)$ using sMLE equations is shown in the right panel of figure 3(a). We see that the basin of clockwise and counterclockwise steady rotating states are smoothly partitioned.
@@
For a larger value of $r=47$, as shown in figure 3(b), we find that sMLE and MLE settles eventually on the steady value of $X$ after a chaotic transient, whereas LE, which does not include hydrodynamic memory, already exhibits permanent chaotic motion. Thus, we already see a signature of delay in the onset of chaotic motion with the inclusion of a memory term in the Lorenz equations. For LE at $Pr=2.5$, the steady rotation branch becomes unstable at $r=32.5$; hence, at $r=47$ the only attracting set in phase space is a strange attractor resulting in chaotic dynamics. We note that the presence of transient chaos for sMLE and MLE indicates a final state sensitivity based on the initial conditions. Unlike $r=3$, where a $Y(0)>0$ would always result in counterclockwise steady rotation and $Y(0)<0$ in clockwise steady rotation, now at $r=47$, whether a clockwise or a counterclockwise steady rotation will be achieved in sMLE and MLE depends sensitively on initial conditions due to transient chaos. This can be seen in the basin of attraction plot of sMLE in the right panel of figure 3(b), where the basins of the two steady states are intricately mixed with fractal boundaries; a characteristic feature of transient chaos [32].
@@
For an even larger value of $r=110$, as shown in figure 3(c), the sMLE and MLE systems have now transitioned to permanent chaos, whereas the LE shows a periodic trajectory. The chaotic nature of the trajectory results in a noisy basin plot for sMLE as shown in the right panel of figure 3(c), where the final velocity at the end of the simulation can now take any value and is not correlated with the initial conditions.
@@
Figure 4: Bifurcation diagram showing the peaks of $|X|$ i.e. $|X_{n}|$ as a function of $r$ and fixed $Pr=2.5$ for (a) standard LE, (b) sMLE with $M(s)=\alpha\,\text{e}^{-s/\alpha}$ and $\alpha=0.5$, and (c) MLE with $\gamma=1$. For each plot, the lower branch is the steady rotation state with the solid curve indicating where it is stable and the dashed curve indicating its unstable. The steady branch becomes unstable at $r=32.5$, $r\approx 73.4$ and $r\approx 176$ for LE, sMLE and MLE, respectively. Note the multistability regions where both chaotic motion and steady rotation coexist. The onset of multistability occurs at $r\approx 25$, $r\approx 51$ and $r\approx 55$ for LE, sMLE and MLE, respectively.
@@
In figure 4 we present a bifurcation diagram showing the peaks of $|X|$ as a function of $r$ and fixed $Pr=2.5$. Here the numerical integration was conducted until a time $s=500$ for the LE and sMLE, and $s=400$ for the MLE (due to the lengthier computations associated with numerical integration of the MLE). For the LE the steady rotation branch becomes unstable at $r=32.5$, but a multistability region where both chaotic motion and steady rotation coexist occurs below this value of $r$. Indeed, chaotic dynamics are first observed at $r\approx 25$. The onset of chaotic dynamics (coexisting with steady rotation) is seen at $r\approx 55$ for the MLE (with $\gamma=1)$, which is a significantly larger $r$ value than for the LE. This larger $r$ value for the second bifurcation in the MLE is also seen in experiments by Peters [20]; that is, the onset of chaos in their experiments occurs at a larger $r$ value than that predicted by the LE, and the present work suggests that hydrodynamic memory is the cause. The sMLE also have an increased $r$ value for the onset of chaotic dynamics (coexisiting with steady rotation) of $r\approx 51$, which suggests this might be a generic consequence of hydrodynamic memory, irrespective of the precise form of the memory kernel.
@@
In figure 5 the multistability of the sMLE (with $\alpha=0.5$, figure 5(a)) and MLE (with $\gamma=1$, figure 5(b)) are examined by plotting the time series of $X$ for two different initial conditions, where one asymptotes towards a chaotic attractor (red) and the other towards a steady state (blue). Figure 5(c) displays basin of attraction plot in the initial condition space $(Y(0),Z(0))$ in the sMLE model for fixed $X(0)=0$ showing the final value of $X$ at the end of simulation time of $400$. We can see coexistence of steady states with smooth separation of their basins, which reside alongside chaotic solutions with noisy values in the basin. Thus, in the multistable region, we see that understanding the size and organization of basin of attraction of steady rotation and chaotic motion might assist with predicting which of the two final states is likely to be realized.
@@
Figure 5: Multistability at $Pr=2.5$ and $r=59$ in the (a) sMLE model with $\alpha=0.5$ and (b) MLE with $\gamma=1$. Time series of $X$ for two different initial conditions where one asymptotes towards a chaotic attractor (red) and the other towards a steady state (blue). (c) Corresponding basin of attraction plot in the initial condition space formed by $(Y(0),Z(0))$ for sMLE model showing the final value of $X$ for fixed $X(0)=0$. Supplemental Videos 4 and 5 show for MLE model the corresponding chaotic motion and steady dynamics, respectively.
@@
In figure 6(a) we present a bifurcation diagram of the MLE showing the peaks of $|X|$ as a function of $r$ at fixed $Pr=2.5$ and $\gamma=0.5$. Here, the onset of chaos is at $r\approx 42$ and there is again a region of multistability in which chaotic and steady motion coexist (with steady rotation stable up to $r\approx 90$). In figure 6(b) and figure 6(c) we plot the times series of $X$ just (b) before ($r=87$) and (c) after ($r=92$), respectively, the end of the stability of the steady rotation branch. At $r=87$ there is a slow convergence of $X$ toward its steady value, and for $r=92$ there is a slow growth of $X$, which we envision would at some time point beyond the simulation become chaotic. The value of $\gamma=0.5$ was chosen as this is approximately the same value as in the experiments of [20]. They observed chaotic dynamics at a field strength of $6.5\mathrm{kV/cm}$, whereas the LE would predict chaos to onset at $5.5\mathrm{kV/cm}$, which corresponds to the ratio of $r$ values (experiment to LE) of around $1.4$. Our computations at $\gamma=0.5$ give a ratio of $r$ values (MLE to LE) of $42/32.5\approx 1.3$ for the onset of chaos. This corresponds to a dimensional onset field strength of $6.3\mathrm{kV/cm}$ for the MLE. Thus the ratio of $r$ values (experiment to MLE) is around $1.06$, which is in encouraging agreement.
To further compare the nature of chaotic dynamics near the onset of chaos between the three models, we plot the first-return map of peaks of $|X|$ in Fig. 7 for LE, sMLE and MLE models at $r=45$ and $Pr=2.5$. As it can be seen in Fig. 7(a), the LE model shows the characteristic cusp-like structure of the first return map that is commonly reported for Lorenz chaos. Alternatively, as shown in Figs. 7(b)-(c), for sMLE and MLE respectively, we find a double-cusp structure in the first return map. This suggests that the double-cusp structure might be a consequence of memory effects in our extended Lorenz sMLE and MLE models. Furthermore, it is not entirely clear whether the cusp structure in the experimental first return map of [20] (see their Fig. 7) has a single cusp or a closely spaced double-cusp structure. This feature might be a useful way to characterize memory effects in chaotic Quincke rotation experiments.
@@
In figure 8 the time series of $X$ is plotted at a large value of $r=250$ for the LE, sMLE with $\alpha=0.5$, and MLE with $\gamma=1$. All three series converge onto a periodic orbit; however, the orbits are not in phase. It is well known for LE that periodic solutions exist at large $r$ and the system becomes integrable in this limit [30]. It appears that we get the same limiting behavior for sMLE and MLE for large $r$. Specifically, we find that for all three models, the period of oscillations decrease whereas the amplitude of oscillations increase, with increasing $r$. Furthermore, we find that the period of the Lorenz-type systems with memory (sMLE and MLE) is larger than that of LE, and the amplitude of the Lorenz-type systems with memory is smaller than that of LE, at a given $r$ in this periodic regime. It would be interesting to explore this regime experimentally and verify if periodic rotations are obtained for Quincke rotors under a large magnitude of applied electric field strengths.
@@
Figure 6: (a) Bifurcation diagram showing the peaks of $|X|$ i.e. $|X_{n}|$ as a function of $r$ and fixed $Pr=2.5$ for full MLE with $\gamma=0.5$. The lower branch is the steady rotation state with the solid curve indicating where it is stable and the dashed curve indicating its unstable. (b) Time series of $X$ just (b) before ($r=87$) and (c) after ($r=92$) the end of the stability of the steady rotation branch.
Figure 7: First return map for the peaks in time series of the magnitude of $X$ for $r=45$ and fixed $Pr=2.5$ for (a) LE, (b) sMLE with $\alpha=0.25$, and (c) full MLE with $\gamma=0.5$.
@@
Figure 8: Time series of $X$ showing periodic oscillations for large $r=250$ and fixed $Pr=2.5$ for (a) LE, (b) sMLE with $\alpha=0.5$, and (c) full MLE with $\gamma=1$. Supplemental Video 6 shows the dynamics of MLE in this periodic state.
VIII Conclusion